# BIV-Priv-Seg: Locating Private Content in Images Taken by People With Visual Impairments

### v1

@@ Abstract
Individuals who are blind or have low vision (BLV) are at a heightened risk of sharing private information if they share photographs they have taken. To facilitate developing technologies that can help preserve privacy, we introduce BIV-Priv-Seg, the first localization dataset originating from people with visual impairments that shows private content. It contains 1,028 images with segmentation annotations for 16 private object categories. We first characterize BIV-Priv-Seg and then evaluate modern models’ performance for locating private content in the dataset. We find modern models struggle most with locating private objects that are not salient, small, and lack text as well as recognizing when private content is absent from an image. We facilitate future extensions by sharing our new dataset with the evaluation server at https://anonymous.com.
@@ 1 Introduction
To facilitate the development of privacy-preserving technologies, we introduce the first localization dataset that originates from people with visual impairments and shows private content. Called BIV-Priv-FewShot, it extends prior work’s [44] collection of images that were taken by BLV photographers of (non-personal) private content. For each of the 1,028 images in the dataset, we collected human-annotated segmentations for every instance of the 16 photographed private object categories. This culminated in $967$ instance segmentations. Examples of annotated images from every private object category are shown in Figure 1.
@@ 3.2 Dataset Analysis
When observing the histogram showing the percentage of images containing different numbers of annotated objects (Figure 3(a)), we observe two key distinctions of our dataset. First, our dataset is the only one with images lacking objects of interest, with this occurring for 18% of the images. These reflect both when no target object is visible and when the target object is a paper positioned on its blank side, rather than the side showing the document’s text (and so private text as well). This amount is similar to that reported by prior work [9], where roughly 20% of BLV photographers’ images lacked the visual evidence needed to answer their visual question despite showing recognizable content [20]. We retain such images to reflect that the photographers intended to capture an object of interest in such images [44], highlighting the challenge of dealing with target-less samples. This is to encourage modern algorithms to achieve a good true negative rate, which is critical for many real-world applications. To support fair dataset comparisons, we exclude these images when analyzing our dataset for the rest of this section. The second key distinction of our dataset is that it rarely has more than one annotated object per image. We attribute this to the distinct photography goal for BIV-Priv [44], where the goal for every image was to photograph a single object, whereas other datasets such as COCO-20i were intentionally curated to show complex scenes showing more than one object (i.e., none-iconic images). When observing the histogram showing the percentage of images containing different numbers segments (i.e., disconnected areas) (Figure 3(b)), we observe that our BIV-Priv-Seg dataset typically exhibits only one segment per object. VizWiz-FewShot, in contrast, more often has the most with up to $122$ segments in an object. While an object itself exists as a unified entity in the real world, multiple counts of segments arise in an object under two circumstances: (1) whenocclusions break objects into more than one region and (2) when objects are partially outside the frame of the image so there are disconnected regions of an object. Our dataset limits these scenarios, thereby permitting more focused model benchmarking analysis for other challenges that can arise, as discussed in this section.
@@ 4.2 VLM Localization
Prompt II. Privacy-aware: We use the following generic prompt “Segment it if there is a private object in the given image, otherwise say there is no private object.”, where the specific category name is replaced with the generic phrase “private object”. This prompt captures is intended to capture model’s ability to understand the concept of privacy.
Prompt III. Privacy-explained: To provide the context of privacy, we provide the list of private categories in our dataset and then request the segmentation as follows: “Private objects include {all category names}. Segment it if there is a private object in the given image, otherwise say there is no private object.” We intentionallydesigned these model prompts toallow for a negative response regarding the presence of the target, meaning that the model does not return a segmentation result. We achieved this by extending the grounding prompts from previous work [39] from“Segment something” to “Segment it if something, otherwise say there is no something”.
@@
Privacy classification: We first evaluate the model performance on two levels: (1) Semantic level prediction, which is based on the language used in the conversation, where positive responses such as “yes, here is the mask” or “sure, here is the mask” indicate the presence of a target, and negative responses such as “no, there is no…” indicate the absence of a target; and (2) Mask level prediction, which is determined by whether any pixel in the generated mask has a positive prediction. The results are presented in accuracy, recall, and true negative rate. Accuracy shows the overall performance; recall reflects the percentage of correct predictions among the target-containing images; true negative rate reflects the percentage of correct predictions among the target-less images. Segmentation: For the images that are classified as private on both semantic and mask levels, we then perform localization. The mIoU and cIoU metrics are used to evaluate the semantic segmentation results across all categories in each scenario. As most of the images in this dataset only contain one target object, we convert those segmentation results into instance segmentation format and evaluate the models in AP50 and Recall metrics. This two-step evaluation highlights the model’s ability to identify whether there are private objects in the image and its localization ability.
@@
We conducted analyses similar to object size and presence of text as we did for few-shot localization algorithms and additional analysis on object saliency. Becausee observed from qualitative results that VLMs tend to segment the salient object in the images, we also evaluated based on whether the target object is the salient object. To do this, we identified whether a target object is salient by applying a salient object detection model [26] that has shown to yield comparable performance to humans for images taken by people with visual impairments (i.e., for the VizWiz dataset) [40] and then applyinga $0.5$ threshol to its intersection over union (IoU) with the target object. This means that when over half of the masks are overlapped, the object is considered salient in our analysis. The results of object size and presence of text align with that of few-shot localization, and we found in the results of object saliency that the model performs better on salient objects.

### v2

@@ Abstract
Individuals who are blind or have low vision (BLV) are at a heightened risk of sharing private information if they share photographs they have taken. To facilitate developing technologies that can help preserve privacy, we introduce BIV-Priv-Seg, the first localization dataset originating from people with visual impairments that shows private content. It contains 1,028 images with segmentation annotations for 16 private object categories. We first characterize BIV-Priv-Seg and then evaluate modern models’ performance for locating private content in the dataset. We find modern models struggle most with locating private objects that are not salient, small, and lack text as well as recognizing when private content is absent from an image. We facilitate future extensions by sharing our new dataset with the evaluation server at https://vizwiz.org/tasks-and-datasets/o
bject-localization/
@@ 1 Introduction
To facilitate the development of privacy-preserving technologies, we introduce the first localization dataset that originates from people with visual impairments and shows private content. Called BIV-Priv-Seg, it extends prior work’s [44] collection of images that were taken by BLV photographers of (non-personal) private content. For each of the 1,028 images in the dataset, we collected human-annotated segmentations for every instance of the 16 photographed private object categories. This culminated in $967$ instance segmentations. Examples of annotated images from every private object category are shown in Figure 1.
@@ 3.2 Dataset Analysis
When observing the histogram showing the percentage of images containing different numbers of annotated objects (Figure 3(a)), we observe two key distinctions of our dataset. First, our dataset is the only one with images lacking objects of interest, with this occurring for 18% of the images. These reflect both when no target object is visible and when the target object is a paper positioned on its blank side, rather than the side showing the document’s text (and so private text as well). This amount is similar to that reported by prior work [9], where roughly 20% of BLV photographers’ images lacked the visual evidence needed to answer their visual question despite showing recognizable content [20]. We retain such images to reflect that the photographers intended to capture an object of interest in such images [44], highlighting the challenge of dealing with target-less samples. This is to encourage modern algorithms to achieve a good true negative rate, which is critical for many real-world applications. To support fair dataset comparisons, we exclude these images when analyzing our dataset for the rest of this section. The second key distinction of our dataset is that it rarely has more than one annotated object per image. We attribute this to the distinct photography goal for BIV-Priv [44], where the goal for every image was to photograph a single object, whereas other datasets such as COCO-20i were intentionally curated to show complex scenes showing more than one object (i.e., none-iconic images).
When observing the histogram showing the percentage of images containing different numbers segments (i.e., disconnected areas) (Figure 3(b)), we observe that our BIV-Priv-Seg dataset typically exhibits only one segment per object. VizWiz-FewShot, in contrast, more often has the most with up to $122$ segments in an object. While an object itself exists as a unified entity in the real world, multiple counts of segments arise in an object under two circumstances: (1) when occlusions break objects into more than one region and (2) when objects are partially outside the frame of the image so there are disconnected regions of an object. Our dataset limits these scenarios, thereby permitting more focused model benchmarking analysis for other challenges that can arise, as discussed in this section.
@@ 4.2 VLM Localization
Prompt II. Privacy-aware: We use the following generic prompt “Segment it if there is a private object in the given image, otherwise say there is no private object.”, where the specific category name is replaced with the generic phrase “private object”. This prompt captures is intended to capture the model’s ability to understand the concept of privacy.
Prompt III. Privacy-explained: To provide the context of privacy, we provide the list of private categories in our dataset and then request the segmentation as follows: “Private objects include {all category names}. Segment it if there is a private object in the given image, otherwise say there is no private object.”
We intentionally designed these model prompts to allow for a negative response regarding the presence of the target, meaning that the model does not return a segmentation result. We achieved this by extending the grounding prompts from previous work [39] from “Segment something” to “Segment it if something, otherwise say there is no something”.
@@
Privacy classification: We first evaluate the model performance on two levels: (1) Semantic level prediction, which is based on the language used in the conversation, where positive responses such as “yes, here is the mask” or “sure, here is the mask” indicate the presence of a target, and negative responses such as “no, there is no…” indicate the absence of a target; and (2) Mask level prediction, which is determined by whether any pixel in the generated mask has a positive prediction. The results are presented in accuracy, recall, and true negative rate. Accuracy shows the overall performance; recall reflects the percentage of correct predictions among the target-containing images; true negative rate reflects the percentage of correct predictions among the target-less images.
Segmentation: For the images that are classified as private on both semantic and mask levels, we then perform localization. The mIoU and cIoU metrics are used to evaluate the semantic segmentation results across all categories in each scenario. As most of the images in this dataset only contain one target object, we convert those segmentation results into instance segmentation format and evaluate the models in AP50 and Recall metrics.
This two-step evaluation highlights the model’s ability to identify whether there are private objects in the image and its localization ability.
@@
We conducted analyses similar to object size and presence of text as we did for few-shot localization algorithms and additional analysis on object saliency. Because we observed from qualitative results that VLMs tend to segment the salient object in the images, we also evaluated based on whether the target object is the salient object. To do this, we identified whether a target object is salient by applying a salient object detection model [26] that has shown to yield comparable performance to humans for images taken by people with visual impairments (i.e., for the VizWiz dataset) [40] and then applying a $0.5$ threshold to its intersection over union (IoU) with the target object. This means that when over half of the masks are overlapped, the object is considered salient in our analysis. The results of object size and presence of text align with that of few-shot localization, and we found in the results of object saliency that the model performs better on salient objects.
@@ 5 Conclusions
Acknowledgments.
This project was supported by a National Science Foundation SaTC award (#2148080, #2126314, and #2125925).